# Geometric Hyperscanning of Affect under Active Inference

### v2

@@ Abstract
Second-person neuroscience holds social cognition as embodied meaning co- regulation through reciprocal interaction, modeled here as coupled active inference with affect emerging as inference over identity-relevant surprise. Each agent maintains a self- model that tracks violations in its predictive coherence while recursively modeling the other. Valence is computed from self-model prediction error, weighted by self- relevance, and modulated by prior affective states and by what we term temporal aiming, which captures affective appraisal over time. This accommodates shifts in the self-other boundary, allowing affect to emerge at individual and dyadic levels. We propose a novel method termed geometric hyperscanning, based on the Forman-Ricci curvature, to empirically operationalize these processes: it tracks topological reconfigurations in inter-brain networks, with its entropy serving as a proxy for affective phase transitions such as rupture, co-regulation, and re-attunement.
Keywords:
Active inference Hyperscanning Second-person neuroscience Dyadic coupling Forman-Ricci curvature Phase transitions.
@@ 1 Introduction
How do agents come to feel ‘together’? How do they navigate the sharing of beliefs about the world and themselves, and the fragile terrain of affective meaning? Traditional models of social cognition often approach this question through the lens of mental state attribution: one agent models another’s beliefs, intentions, and emotions. This framing which we call adopting a third-person stance towards someone, has dominated both theoretical and empirical social cognition research [9]. However, such models ignore a crucial fact: social understanding, in most of its ecologically valid forms, does not unfold in detached observation but in active, embodied engagement with others. As Schilbach et al. [37] argue, social interaction is not simply the backdrop against which inference occurs; instead, it is itself the medium through which meaning is constituted. This shift, often captured under the rubric of ‘second-person neuroscience’ [35, 38], reframes the unit of analysis from the isolated agent to the dyad as a generative system in its own right, by taking mutual engagement in social interaction, rather than social observation, as its central explanandum. Recent formalizations of this framework through the lens of active inference [29] have provided the missing computational scaffolding: they model interacting agents as mutually coupled generative systems that co-regulate meaning via recursive belief updating. Social understanding, in this view, is not computed about another agent, but emerges while in social interaction with another through ongoing cycles of expectation, its violation, and realignment. The dyad becomes a unit of analysis in its own right and can be regarded as a shared generative manifold, which can be modeled as a coupled system of intra- and interpersonal processes across multiple modalities [5, 6].
@@ 1.1 The Affective Gap
We extend second-person active inference to the domain of affect, modeling it as recursive inference over two types of prediction error: mismatch between predictions and outcomes relevant to the world, and those relevant to identity. Drawing on Jiang and Luo’s [26, 25] model, valence reflects an inference over the integrity of the self-model, embedded in the dyadic generative manifold. Because both agents maintain self-models while recursively modeling each other, affective dynamics become entangled, shaping belief updating and action selection. This recursive process promotes interpersonal attunement and stabilization [6, 42, 39], while affective rupture triggers cognitively costly recalibration [10, 3]. We propose geometric hyperscanning to link these formal dynamics to neural signatures, expanding the framework of second-person active inference and its formal dynamics of belief and affect to empiric research.
This paper makes three contributions. First, we formalize affect as recursive inference over self-model coherence, modeling valence as identity-relevant prediction error weighted by self-relevance, modulated by affective memory, and shaped by temporal aiming — the agent’s orientation across past and future affective states (Section 2.1). We show how recursive modeling of the other dynamically shifts self–other boundaries (Section 2.2). Second, we introduce geometric hyperscanning, an empirical method based on Forman-Ricci curvature (FRc), to track topological reconfigurations in inter-brain networks and infer affective phase transitions (Section 2.3). Third, we integrate this formal-empirical framework within second-person neuroscience and active inference, offering a scalable architecture for modeling recursive affective dynamics across psychotherapy, development, and naturalistic interaction, and outline its broader implications (Section 3).
@@ 2 Second-Person Neuroscience and Dyadic Active Inference
The recursive nature of this system (wherein both agents model each other modeling them) does not lead to an infinite regress. Rather, as Friston and Frith (2015) demonstrated, it dissolves into a shared dynamic narrative, where both agents come to predict and enact the same process. In other words, both agents use the same model and the dyad, thus, becomes the minimal unit of inference.
In contrast, traditional third-person approaches to social cognition, such as Theory of Mind or mindreading paradigms cast social understanding as an observer inferring another’s hidden mental states. Classic experiments, for instance, might have a participant watch a video or read a story and then guess what another person believes or intends. These paradigms, while valid, involve a one-way attribution process, often in non-interactive settings. Neuroscientifically, third-person “mentalizing” tasks reliably engage a network including the dorsomedial prefrontal cortex (dmPFC), ventromedial prefrontal cortex (vmPFC), and temporoparietal junction (TPJ), which is often called the mentalizing or theory-of-mind network [19]. But in a live interaction, one’s brain is not only modeling the other but also responding to them in real time, creating a feedback loop. This form of social connectedness has measurable two-brain dynamics: an engaged “I–Thou” exchange produces alignment of neural rhythms that does not occur in a one-directional “I observe him” situation (see [14, 38] for a recent review).
[20] refers to these coupled networks as mutual social attention systems, proposing that when two people directly attend to each other, their brains temporarily form an integrated system. The second-person emphasis on real-time mutual engagement dovetails with active inference, because both paradigms recognize that understanding others is an active, dynamical, and bidirectional process of alignment.
@@ 2.1 Valence as Inference over the Self-Model
Within this coupled generative architecture, affect is conceptualized as dynamic regulatory processes guiding cognition and behavior on the basis of the integrity and coherence of an agent’s predictive self-model [26, 25]. Central to this approach is the idea that emotional valence signals the alignment or mismatch between predicted and observed outcomes, particularly those relevant to identity and social expectations [3].
@@
Figure 1: Temporal aiming in affective inference. The figure illustrates how agents A evaluate self-model prediction error over time according to different temporal orientations. Affect is shaped not only by the magnitude and relevance of prediction error, but also by the agent’s temporal aim, i.e. whether inference is directed toward anticipated futures or prior experiences, and whether updates occur rapidly or gradually. This temporal structuring of affect modulates both the stability and flexibility of policy selection within dyadic interaction.
This modulation adds an important dimension to affective inference: agents may respond differently to the same interaction depending on whether they are prospectively anxious, retrospectively regretful, or temporally stable. Temporal aiming thus complements self-relevance and prediction error as core variables in our model of valence. We conceptualize emotions as functions of valence (alignment of identity-relevant predictions) and arousal (uncertainty and intensity of self-relevant evaluations), with both dimensions anchored in the evolving self-concept, shaped by reflection, comparison, and internalized ideals. In this framework, valence reflects the agent’s appraisal of ongoing interaction, computed as a weighted function of self-model prediction error and self-relevance, and modulated by prior affective memory [25]:
@@
  • $PE_{\text{self},t}$ is the prediction error between expected and actual self-state,
  • $SR_{t}$ is a dynamic scalar indexing the self-relevance of the context,
  • $V_{SR,t}$ is retrieved from similar prior contexts,
  • $\alpha,\beta$ are precision-modulating weights.
It follows that whenever the agent experiences a self-related prediction error, it generates an affective response proportional to that error and to how self-relevant it is. The parameter $\alpha$ scales the impact of immediate prediction error on valence. An agent $i$ with a larger $\alpha$ will exhibit heightened emotional sensitivity to surprises—particularly negative surprises, which yield more negative valence. The second term, $\beta\cdot V_{SR,t}$ with $10<\beta<$ 1, makes valence a leaky integrator: it carries forward some fraction of previous valence (mood) into the current moment. The term $PE_{\text{self},t}$ includes both the agent’s previous valence $V(t-1)$ and contributions from past episodes that had similar self-relevance or context as the current one. We can treat $V_{SR,t}$ as an exponentially decaying sum of past valences, with greater weight assigned to more recent times and to events with comparable self-relevance (SR). In the notation of the original model, self-model prediction error and self-relevance map onto $PE_{\text{self}}$ and $SR$, respectively, where the last term of $V_{SR,t}$ helps to maintain the affective experience as iterative and continuous.
@@ 2.1.1 Valence under the Dyadic Inference Framework
In the dyadic active inference framework, each agent’s self-model is treated as a dynamic function of its internal world model, which includes a simulated model of the other agent and incoming interactional feedback. At time $t$, agent $i$’s self-model is conditioned on beliefs about the other agent’s state, actions, and observations:
@@
Agent $i$’s full generative model factorizes as:
@@
Within this framework, valence ($V_{i},t$) reflects the agent’s affective appraisal of self- model coherence over time. It is computed from the prediction error in the self-model, modulated by both $SR_{t}$ and $V_{SR,t}$. The expected valence is defined as:
@@
This expresses the agent’s belief about the likelihood that future outcomes will sustain its internalized social narrative, given its evolving model of the partner and ongoing interactional feedback.
@@ 2.1.2 Affective Valence as a Prior Over Policy in Planning-as-Inference
By design, valence serves as an internal reward/safety signal: high positive valence indicates smooth, effective interaction and affective synchrony, while negative valence signals that current synchrony is underperforming: prompting either a new attempt to restore alignment or a withdrawal from the interaction, depending on individual differences in valence construction and perception (modeled by the parameters $\alpha$ and $\beta$). Though we do not explicitly model all downstream effects here, valence functions as a feedback loop, providing a readout of each agent2̆019s internal model performance and modulating subsequent perception and action. In our model, affect modulates the posterior over policies, effectively serving as a soft prior that biases action selection toward more affectively coherent trajectories. Following Da Costa’s (2020) [13] planning-as-inference paradigm, we express policy selection as:
@@
  • $\lambda$ is an inverse temperature parameter controlling affective precision or confidence in the valence signal, and
  • $g(V_{t},\pi_{t})$ is a compatibility function mapping affective coherence to policy likelihood.
This captures the recursive and value-sensitive nature of planning in social contexts. The agent is more likely to choose plans that are consistent with positive valence (i.e., minimal surprise and maximal narrative alignment), while negative valence reduces the likelihood of maintaining the current trajectory. When both agents are engaged in this recursive affective inference, the system forms a dynamically coupled manifold —a shared generative space in which each agent’s self- coherence is entangled with the other’s behavior. Moments of affective rupture mark local minima in joint narrative alignment and correspond to peaks in dyadic expected free energy. Repairing these ruptures requires coordinated adjustments in beliefs and policies across both agents. This framework completes a recursive loop: expected valence, derived from self-model integrity and anticipated partner behavior, acts as a prior over policy selection, guiding the agent toward actions expected to preserve or restore affective coherence. These chosen policies then lead to new interactions and observations, which update the agent’s self-model and generate realized valence. In turn, the mismatch between expected and realized valence yields an affective prediction error that shapes the next cycle of inference. This recursive loop (Fig. 2) summarizes how affectively modulated policies guide dyadic interaction through cycles of prediction error, self-model update, and realignment.
@@ 2.2 Recursive Coupling and Shared Surprise
In exploring the dynamics of affective interactions, it is useful to draw upon the concept of intra-action from neo-materialist philosophy [4]. Intra-action emphasizes that interacting entities do not preexist independently but rather emerge through their mutual entanglement. Under this lens, the dyadic interaction is not simply a transfer of emotional states between two separate agents; instead, it represents a continual co-construction of affective experiences and identities [33]. Affective states, from this perspective, are inherently recursive and relational, arising from the iterative feedback loops between agents. Affective ruptures, moments of misalignment or increased prediction error, serve as significant informational signals indicating points of divergence in the shared expectations of interacting agents. These affective evaluations modulate the posterior over policies, biasing action selection toward trajectories that restore narrative alignment and reduce identity-relevant surprise. Such an entanglement of affect and policy selection turns the dyad into a dynamic field of co-regulated affective inference, where both agents are simultaneously minimizing joint expected free energy:
@@
At the dyadic level, recursive coupling of valence creates a shared attractor structure: higher valence biases policy selection toward trajectories that preserve narrative coherence and reduce expected future free energy. Conversely, negative valence can trigger exploratory policies aimed at restoring alignment or withdrawing from destabilizing interactional trajectories. In this way, valence modulates both immediate affective dynamics and the agent’s future epistemic and instrumental uncertainty.

### v3

@@ Abstract
Second-person neuroscience holds social cognition as embodied meaning co-regulation through reciprocal interaction, modeled here as coupled active inference with affect emerging as inference over identity-relevant surprise. Each agent maintains a self-model that tracks violations in its predictive coherence while recursively modeling the other. Valence is computed from self-model prediction error, weighted by self-relevance, and modulated by prior affective states and by what we term temporal aiming, which captures affective appraisal over time. This accommodates shifts in the self-other boundary, allowing affect to emerge at individual and dyadic levels. We propose a novel method termed geometric hyperscanning, based on the Forman-Ricci curvature, to operationalize these processes empirically: it tracks topological reconfigurations in inter-brain networks, with its entropy serving as a proxy for affective phase transitions such as rupture, co-regulation, and re-attunement.
@@ 1 Introduction
How do agents come to feel ‘together’? How do they navigate the sharing of beliefs about the world and themselves, and the fragile terrain of affective meaning? Traditional models of social cognition often approach this question through the lens of mental state attribution: one agent models another’s beliefs, intentions, and emotions. Adopting a third-person stance towards someone has dominated theoretical and empirical social cognition research [9]. However, such models ignore a crucial fact: social understanding, in most of its ecologically valid forms, does not unfold in detached observation but in active, embodied engagement with others. As Schilbach et al. [37] argue, social interaction is not simply the backdrop against which inference occurs; instead, it is the medium through which meaning is constituted. This shift, often captured under the rubric of ‘second-person neuroscience’ [35, 38], reframes the unit of analysis from the isolated agent to the dyad as a generative system in its own right, by taking mutual engagement in social interaction, rather than social observation, as its central explanandum. Recent formalizations of this framework through the lens of active inference [29] have provided the missing computational scaffolding: they model interacting agents as mutually coupled generative systems that co-regulate meaning via recursive belief updating. In this view, social understanding is not computed about another agent, but emerges while in social interaction with another through ongoing cycles of expectation, violation, and realignment. The dyad becomes a unit of analysis in its own right and can be regarded as a shared generative manifold, which can be modeled as a coupled system of intra- and interpersonal processes across multiple modalities [5, 6].
@@ 1.1 The Affective Gap
We extend second-person active inference to the domain of affect, modeling it as recursive inference over two types of prediction error: mismatch between predictions and outcomes relevant to the world, and those relevant to identity. Drawing on Jiang and Luo’s [26, 25] model, valence reflects an inference over the integrity of the self-model, embedded in the dyadic generative manifold. Because both agents maintain self-models while recursively modeling each other, affective dynamics become entangled, shaping belief updating and action selection. This recursive process promotes interpersonal attunement and stabilization [6, 42, 39], while affective rupture triggers cognitively costly recalibration [10, 3]. We propose geometric hyperscanning to link these formal dynamics to neural signatures, expanding the framework of second-person active inference and its formal dynamics of belief and affect to empirical research.
This paper makes three contributions. First, we formalize affect as recursive inference over self-model coherence. We model valence as identity-relevant prediction error weighted by self-relevance, modulated by affective memory, and shaped by temporal aiming — the agent’s orientation across past and future affective states (Section 2.1). We show how recursive modeling of the other dynamically shifts self–other boundaries (Section 2.2). Second, we introduce geometric hyperscanning, an empirical method based on Forman-Ricci curvature (FRc), to track topological reconfigurations in inter-brain networks and infer affective phase transitions (Section 2.3). Third, we integrate this formal-empirical framework within second-person neuroscience and active inference. We offer a scalable architecture for modeling recursive affective dynamics across psychotherapy, development, and naturalistic interaction, and outline its broader implications (Section 3).
@@ 2 Second-Person Neuroscience and Dyadic Active Inference
The recursive nature of this system (wherein both agents model each other, modeling them) does not lead to an infinite regress. Instead, as Friston and Frith (2015) demonstrated, it dissolves into a shared dynamic narrative, where both agents come to predict and enact the same process. In other words, both agents use the same model and the dyad, thus, becomes the minimal unit of inference.
In contrast, traditional third-person approaches to social cognition, such as Theory of Mind or mindreading paradigms, cast social understanding as an observer inferring another’s hidden mental states. Classic experiments, for instance, might have a participant watch a video or read a story and then guess what another person believes or intends. These paradigms, while valid, involve a one-way attribution process, often in non-interactive settings. Neuroscientifically, third-person “mentalizing” tasks reliably engage a network including the dorsomedial prefrontal cortex (dmPFC), ventromedial prefrontal cortex (vmPFC), and temporoparietal junction (TPJ), which is often called the mentalizing or theory-of-mind network [19]. However, in a live interaction, one’s brain is modeling the other and responding to them in real time, creating a feedback loop. This form of social connectedness has measurable two-brain dynamics: an engaged “I–Thou” exchange produces alignment of neural rhythms that does not occur in a one-directional “I observe him” situation (see [14, 38] for a recent review).
[20] refers to these coupled networks as mutual social attention systems, proposing that when two people directly attend to each other, their brains temporarily form an integrated system. The second-person emphasis on real-time mutual engagement dovetails with active inference, because both paradigms recognize that understanding others is an active, dynamical, and bidirectional alignment process.
@@ 2.1 Valence as Inference over the Self-Model
Within this coupled generative architecture, affect is conceptualized as dynamic regulatory processes guiding cognition and behavior based on the integrity and coherence of an agent’s predictive self-model [26, 25]. Central to this approach is the idea that emotional valence signals the alignment or mismatch between predicted and observed outcomes, particularly those relevant to identity and social expectations [3].
@@
Figure 1: Temporal aiming in affective inference. The figure illustrates how agents evaluate self-model prediction error over time according to different temporal orientations. Affect is shaped not only by the magnitude and relevance of prediction error, but also by the agent’s temporal aim, i.e., whether inference is directed toward anticipated futures or prior experiences, and whether updates occur rapidly or gradually. This temporal structuring of affect modulates the stability and flexibility of policy selection within dyadic interaction.
This modulation adds an essential dimension to affective inference: agents may respond differently to the same interaction depending on whether they are prospectively anxious, retrospectively regretful, or temporally stable. Temporal aiming thus complements self-relevance and prediction error as core variables in our valence model. We conceptualize emotions as functions of valence (alignment of identity-relevant predictions) and arousal (uncertainty and intensity of self-relevant evaluations), with both dimensions anchored in the evolving self-concept, shaped by reflection, comparison, and internalized ideals. In this framework, valence reflects the agent’s appraisal of ongoing interaction or how “good” or “bad” the current moment feels to the agent, computed as a weighted function of self-model prediction error and self-relevance, and modulated by prior affective memory [25]:
@@
  • $PE_{\text{self},t}$ is the prediction error between expected and actual self-state, capturing the “surprise” the agent experiences about itself
  • $SR_{t}$ is a dynamic scalar indexing the self-relevance of the context, reflecting how much the situation matters for the agent’s identity
  • $V_{SR,t}$ is retrieved from similar prior contexts, representing the lingering emotional tone (mood) shaped by related past experiences,
  • $\alpha,\beta$ are precision-modulating weights, that determine how strongly the agent reacts to immediate surprises versus how much past mood carries over..
It follows that whenever the agent experiences a self-related prediction error, it generates an affective response proportional to that error and to how self-relevant it is. The first term, $\alpha\cdot PE_{\text{self},t}\cdot SR_{t}$, encodes the instantaneous emotional impact: large, self-relevant surprises cause large valence shifts. The parameter $\alpha$ scales the impact of immediate prediction error on valence. An agent $i$ with a larger $\alpha$ will exhibit heightened emotional sensitivity to surprises, particularly negative ones, which yield more negative valence. The second term, $\beta\cdot V_{SR,t}$ with $0<\beta<1$, he second term, $\beta\cdot V_{SR,t}$, introduces emotional inertia: it blends in past affective states, so that current mood is partly a continuation of recent emotional history, it makes valence a leaky integrator: it carries forward some fraction of previous valence (mood) into the current moment. The term $PE_{\text{self},t}$ includes both the agent’s previous valence V(t1)V(t-1) and contributions from past episodes that had similar self-relevance or context as the current one. We can treat $V_{SR,t}$ as an exponentially decaying sum of past valences, with greater weight assigned to more recent times and to events with comparable self-relevance (SR). In the notation of the original model, self-model prediction error and self-relevance map onto $PE_{\text{self}}$ and $SR$, respectively, where the last term of $V_{SR,t}$ helps to maintain the affective experience as iterative and continuous.
@@ 2.1.1 Valence under the Dyadic Inference Framework
In dyadic active inference, each agent’s self-model is treated as a dynamic function of its internal world model, which includes a simulated model of the other agent and incoming interactional feedback. At time $t$, agent $i$’s self-model is conditioned on beliefs about the other agent’s state, actions, and observations:
@@
This reflects that one’s self-concept in a social setting is co-constructed by self-generated and socially observed information. Agent $i$’s full generative model factorizes as:
@@
This expresses that (1) the agent’s observations depend on its state and actions, (2) its actions depend on both its own and the other’s state, and (3) its own state is partly inferred from the other’s state. Valence ($V_{i},t$) measures how emotionally coherent the agent’s self-model feels over time. It is computed from the prediction error in the self-model, modulated by both $SR_{t}$ and $V_{SR,t}$. The expected valence is defined as:
@@
This means that valence is the expectation, over possible outcomes, of how well those outcomes will sustain the agent’s internalized social narrative, given its evolving model of the partner and the feedback from ongoing interaction.
@@ 2.1.2 Affective Valence as a Prior Over Policy in Planning-as-Inference
By design, valence serves as an internal reward/safety signal: high positive valence indicates that the interaction is proceeding smoothly, with strong affective synchrony, while negative valence signals a breakdown in synchrony, prompting either a new attempt to restore alignment or a withdrawal from the interaction, depending on individual differences in valence construction and perception (modeled by the parameters $\alpha$ and $\beta$). Which path the agent takes depends on individual differences in emotional sensitivity and persistence, captured by the parameters $\alpha$ and $\beta$ Though we do not explicitly model all downstream effects here, valence functions as a feedback loop, a readout of the self-model’s performance and modulating subsequent perception and action. In practical terms, this means that affect shapes the space of possible future actions by making certain policies feel “more natural” or “safer” than others. In our model, affect modulates the posterior over policies, effectively serving as a soft prior that biases action selection toward more affectively coherent trajectories. Following Da Costa’s (2020) [13] planning-as-inference paradigm, we express policy selection as:
@@
  • $\lambda$ is an inverse temperature parameter controlling affective precision or confidence in the valence signal (high $\lambda$ = more deterministic choice; low $\lambda$ = more exploratory choice), and
  • $g(V_{t},\pi_{t})$ is a compatibility function mapping affective coherence to policy likelihood (higher compatibility = higher policy probability).
This captures the recursive and value-sensitive nature of planning in social contexts. The agent is more likely to choose plans consistent with positive valence (i.e., minimal surprise and maximal narrative alignment), while negative valence reduces the likelihood of maintaining the current trajectory. When both agents are engaged in this recursive affective inference, the system forms a dynamically coupled manifold —a shared generative space in which each agent’s self-coherence is entangled with the other’s behavior. Moments of affective rupture mark local minima in joint narrative alignment and correspond to dyadic expected free energy peaks. Repairing these ruptures requires coordinated adjustments in beliefs and policies across both agents. This framework completes a recursive loop: expected valence, derived from self-model integrity and anticipated partner behavior, acts as a prior over policy selection, guiding the agent toward actions expected to preserve or restore affective coherence. These chosen policies lead to new interactions and observations, which update the agent’s self-model and generate realized valence. In turn, the mismatch between expected and realized valence yields an affective prediction error that shapes the next inference cycle. This recursive loop (Fig. 2) summarizes how affectively modulated policies guide dyadic interaction through cycles of prediction error, self-model update, and realignment.
@@ 2.2 Recursive Coupling and Shared Surprise
In exploring the dynamics of affective interactions, it is helpful to draw upon the concept of intra-action from neo-materialist philosophy [4]. Intra-action emphasizes that interacting entities do not preexist independently but emerge through mutual entanglement. Under this lens, the dyadic interaction is not simply a transfer of emotional states between two separate agents; instead, it represents a continual co-construction of affective experiences and identities [33]. From this perspective, affective states are inherently recursive and relational, arising from the iterative feedback loops between agents. Affective ruptures, moments of misalignment or increased prediction error, serve as significant informational signals indicating points of divergence in the shared expectations of interacting agents. These affective evaluations modulate the posterior over policies, biasing action selection toward trajectories that restore narrative alignment and reduce identity-relevant surprise. Such an entanglement of affect and policy selection turns the dyad into a dynamic field of co-regulated affective inference, where both agents are simultaneously minimizing joint expected free energy:
@@
Here, $G_{\text{joint}}$ quantifies the shared divergence between the dyad’s joint beliefs and its joint generative model. High joint valence corresponds to low joint free energy (smooth alignment), whereas affective ruptures correspond to spikes in joint free energy (misalignment and uncertainty). At the dyadic level, recursive coupling of valence creates a shared attractor structure: higher valence biases policy selection toward trajectories that preserve narrative coherence and reduce expected future free energy. Conversely, negative valence can trigger exploratory policies to restore alignment or withdraw from destabilizing interactional trajectories. This way, valence modulates immediate affective dynamics and the agent’s future epistemic and instrumental uncertainty.